Choctaw Co. completed the following transactions in Year 1, the first year of operation.

1. Issued 35,000 shares of $10 par common stock for $10 per share.

2. Issued 4,500 shares of $20 stated value preferred stock for $20 per share.

3. Purchased 2,500 shares of common stock as treasury stock for $12 per share.

4. Declared a $3,500 cash dividend on preferred stock.

5. Sold 1,200 shares of treasury stock for $14 per share.

6. Paid $3,500 cash for the preferred dividend declared in Event 4.

7. Earned cash revenues of $108,000 and incurred cash expenses of $56,000.

8. Appropriated $9,500 of retained earnings.

Required:

a. Organize the transaction in accounts under an accounting equation.

b. Prepare a balance sheet as of December 31, Year 1.  Required A Required B  Complete this question by entering your answers in the tabs below. Organize the transaction in accounts under an accounting equation. (Enter any decreases to account balances with a minus sign. Select "NA" if there is no effect on the "A Earnings".)

POST IN EXCEL SPREADSHEET Please.
Answer #1

A.

Choctaw Co.
Accounting Equation for the Year 1
Event Assets = Liabilities Stockholders' Equity Accounting Title Retained Earnings
Cash = Dividends
Payable
+ Preferred
Stock
+ Common Stock + Paid-in Capital In Excess Treasury Stock - Treasury Stock + Retained Earnings + Appropriated Retained Earnings
1 350000 = + + 350000 + - + + NA
2 90000 = + 90000 + + - + + NA
3 -30000 = + + + - 30000 + + NA
4 = 3500 + + + - + -3500 + Dividends
5 16800 = + + + 2400 - -14400 + + NA
6 -3500 = -3500 + + + - + + NA
7a. 108000 = + + + - + 108000 + Revenues
7b. -56000 = + + + - + -56000 + Expenses
8 = + + + - + -9500 + 9500 Appropriated retained earnings
Totals 475300 = 0 + 90000 + 350000 + 2400 - 15600 + 39000 + 9500

B.

CHOCTAW CO.
As of December 31, Year 1
Stockholders' equity
Preferred stock 90000
Common stock 350000
Paid-in Capital Treasury stock 2400
Total Paid-In Capital 442400
Retained Earnings
Unappropriated 39000
Appropriated 9500
Total Retained Earnings 48500
Less: Treasury stock -15600
Total Stockholders' Equity 475300